# TOMG-Bench: Evaluating LLMs on Text-based Open Molecule Generation

### v1

@@ 3.3 Metric Design
For the MolCustom task, which includes subtasks such as AtomNum, BondNum, and FunctionalGroup, the following metrics are employed:
  • Accuracy: This metric measures the proportion of generated molecules that successfully meet the criteria. A high accuracy indicates that the LLM is effectively generating molecules according to the request.
  • Novelty: Novelty is a critical metric for MolCustom tasks, as it assesses the ability of LLMs to discover new molecular structures. This is particularly important for potential drug or material candidates, where innovation can lead to significant advancements.
  • Validity: This metric ensures that the generated molecules are chemically valid and follow the rules of molecular syntax.
@@
For the MolEdit and MolOpt tasks, the following metrics are adopted:
  • Accuracy: Similar to the MolCustom category, accuracy here measures the proportion of edited or optimized molecules that could pass the test process.
  • Similarity: Similarity between the original molecule and the edited molecule is an important point to concern. Since there are multiple valid solutions, only considering the accuracy is not sufficient for verification. For implementation, a high similarity indicates that the generated molecule can be modified from the original molecule via as few as possible steps. For instance, when removing a specific functional group, the edited molecule should still be similar to the original molecule. This process should not transform the molecule into an entirely new one nor inadvertently remove additional components.
  • Validity: Similarly, validity is still a crucial metric to confirm that the edited or optimized molecules adhere to the principles of the molecule textual representation.
Notably, the calculation of novelty and similarity metrics only considers valid molecules. Specifically, for the novelty metric, we assess the similarity between the generated molecule and those within the Zinc-250K database.
@@
To comprehensively evaluate the average performance of LLMs on TOMG-Bench, we introduce a weighted average accuracy to rank the performance of LLMs. Considering that the novelty scores for MolCustom and similarity scores for MolEdit and MolOpt are also crucial metrics for evaluating the performance, especially the similarity scores, which help identify correct molecule editing operations. In this case, we adopt the novelty and similarity scores as the weights, and the weighted average accuracy can be computed as follows:
$\bar{w\!Acc}\!=\!\frac{1}{9}\!(\sum_{t\in\!\!M\!o\!l\!C\!u\!s\!t\!o\!m\!}\!(\!n_{t}*Acc_{t}\!)\!+\sum_{t\in\!\{\!M\!o\!l\!E\!d\!i\!t,\!M\!o\!l\!O\!p\!t\!\}}\!(\!s_{t}*Acc_{t}\!)),$ (1)
@@
where $n_{t}$ represents the novelty score for the MolCustom tasks and $s_{t}$ is the similarity score for the MolEdit and MolOpt tasks. $Acc_{t}$ is the accuracy for each subtask $t$. This weighted average accuracy, $\bar{wAcc}$, provides a balanced measure of performance that considers both the accuracy and quality of the generation.
@@ 3.4 OpenMolIns: Instruction Tuning Dataset
We collect the instruction tuning dataset by introducing the samples of the nine subtasks in equal amounts. We still apply the RDKit toolbox to construct the instruction tuning dataset.
In the MolCustom domain, constructing the instruction tuning samples is rather straightforward. We calculate the molecular statistics and encapsulate them within prompts. For example, for AtomNum, we count all the atoms we would consider in the molecule, including their types and numbers. Then, we wrap these statistics with a random pre-defined prompt to construct the training sample. This method allows the generated molecules to better fit the distribution of molecular space.
For the MolEdit and MolOpt domains, we randomly select a functional group from the original molecule for addition or removal, utilizing the RDKit toolbox to execute these operations. Then, similarly, we wrap the original molecule and the edited molecule with a random pre-defined prompt to construct the training sample for MolEdit. In the case of MolOpt, we also evaluate the direction of the desired property changes by the functions in RDKit to determine whether the operation improves or decreases the property value, which is then altogether wrapped by the pre-defined prompts.
To investigate the impact of data scales on the performance of LLMs, we have established five distinct data levels tailored for different training purposes: light, small, medium, large, and xlarge, as illustrated in Section 3.5. Each level represents a different quantity of data, shown in Table 1, allowing us to analyze how the amount of training data influences the model’s ability to learn and generate or edit molecules effectively.
Item Data Size
TOMG-Bench
subtask 5,000
OpenMolIns
light 4,500
small 18,000
medium 45,000
large 90,000
xlarge 1,200,000
@@ 3.5 Statistics
In this section, we outline the basic statistics of TOMG-Bench along with the OpenMolIns dataset. Table 1 shows the details of the data size in TOMG-Bench as well as the OpenMolIns datset. For TOMG-Bench, we have three main tasks with nine subtasks in total, where each subtask contains 5,000 test samples. More details can be found in Appendix A.
For OpenMolIns, we have five distinct data scales: light, small, medium, large, and xlarge, ranging from 4,500 to 1,200,000 examples, which helps us investigate the data scaling law of applying LLMs to the Text-based Open Molecule Generation task. Notably, the nine subtasks within the TOMG-Bench are uniformly distributed in the OpenMolIns dataset.
@@ 4 Experiments
In this section, we present the experiment setup and results. Then, we illustrate our findings based on the observations.
4.1 Setup
4.1.1 Models
@@
Proprietary Models. This category includes LLMs that are only accessible via commercial API services. In this work, we benchmark GPT-4o, GPT-4-turbo, GPT-3.5-turbo Achiam et al. (2023), Claude-3.5 Anthropic (2024b), Claude-3 Anthropic (2024a), and Gemini-1.5-pro Deepmind (2024).
Open-source General LLMs. This group contains open-source LLMs that are tuned with the instruction following capability, which can be used for a wide range of tasks and applications. Specifically, we benchmark Llama-3-70B-Instruct, Llama-3-8B-Instruct, Llama-3.1-8B-Instruct, Llama-3.2-1B-Instruct Dubey et al. (2024), Mistral-7B-Instruct-v0.2 Jiang et al. (2023), Qwen2-7B-Instruct Yang et al. (2024), yi-1.5-9B Young et al. (2024), and chatglm-9B GLM et al. (2024).
Open-source ChEBI-20 Fine-tuned LLMs. LLMs fine-tuned on the ChEBI-20 dataset can grasp some extent of text-based molecule generation capability. In this case, our experiments also cover LLMs like MolT5-small, MolT5-base, MolT5-large Edwards et al. (2022), and BioT5-base Pei et al. (2023).
OpenMolIns Fine-tuned LLMs We further adopt LLMs like Galactica-125M Taylor et al. (2022), Llama3.2-1B-Instruct, and Llama-3.1-8B-Instruct on OpenMolIns dataset for instruction tuning. We specifically include the experiments on five distinct data sizes of OpenMolIns for Galactica-125M.
4.1.2 Implementation Details
We implement various scripts to facilitate the testing of the aforementioned models. For proprietary models, we adopt the OpenAI API  [^1] https://platform.openai.com/docs/ framework. For open-source general LLMs, we utilize both the VLLM  [^2] https://docs.vllm.ai framework and the OpenAI framework. For the remaining LLMs, we adopt the Hugging Face transformers library  [^3] https://huggingface.co/docs/transformers/ for inference. Detailed hyper-parameters are demonstrated in Appendix B.
Furthermore, it is important to note that BioT5 is designed to use SELFIES as input instead of SMILES. Consequently, we convert the molecule SMILES strings into SELFIES format on BioT5.
4.2 Results
Figure 3: The performance of models benchmarked in TOMG-Bench. In TOMG-Bench, LLMs are divided into 4 categories: Proprietary Models, Open-source General LLMs, Open-source ChEBI-20 Fine-tuned LLMs, and OpenMolIns Fine-tuned LLMs. Models whose parameters are known are plotted as dots, while models of unknown parameters are denoted as horizontal lines.
Figure 3 presents the performances and model sizes of different models benchmarked on TOMG-Bench, as well as the instruction-tuning performance of Galactica-125M on the five distinct data levels of OpenMolIns. More precise details are further illustrated in Appendix C.
@@ 4.3 Findings
Based on the above results, we observe the following key findings:
Text-based open molecule generation is challenging for LLMs. As illustrated in Figure 3, we have calculated the weighted average accuracy across all the nine subtasks. Among the LLMs benchmarked, Claude-3.5 stands out as the top performer, achieving a weighted average accuracy of 35.92%. Gemini-1.5-pro follows closely with a weighted average accuracy of 34.80%. These results underscore the considerable scope for improvement, even among the most advanced proprietary LLMs.
It is also worth noting that while the most advanced LLMs like Claude-3.5 and GPT-4o exhibit relatively strong performance in the MolEdit and MolOpt tasks, the more challenging MolCustom task still remains a challenge. In MolCustom tasks, no LLM has managed to achieve an accuracy exceeding 25% for a single subtask. This observation indicates that the generation of molecules from scratch demands a deep understanding of the molecular structural space, an area where current models are still striving to make significant strides.
Most powerful open-source general LLMs can already outperform GPT-3.5-turbo. In the TOMG-Bench, Llama-3-70B-Instruct has achieved an impressive weighted average accuracy of 23.93%, notably outperforming GPT-3.5-turbo, which scored 18.58%. Despite previously lagging behind proprietary models, open-source general LLMs have rapidly bridged the gap. The evolution of the Llama series, in particular, has been remarkable, finally surpassing GPT-3.5-turbo and demonstrating the fast development of open-source general LLMs.
More powerful LLMs inherit a better performance in TOMG-Bench. Across all the LLMs we benchmarked, a clear trend emerged: the more powerful the LLM is, the higher performance it can achieve on the TOMG-Bench. For instance, the GPT series has consistently demonstrated improved performance with each new iteration from GPT-3.5-turbo to GPT-4o.
Similarly, within the Llama-3 series, we could also observe that larger models tend to achieve superior results on the TOMG-Bench. These findings underscore a strong correlation between an LLM’s capabilities and its performance in our benchmark.
@@
ChEBI-20 dataset is insufficient for LLMs to master molecular structures and editing operations. The ChEBI-20 dataset and the associated molecule-caption translation task are designed to bridge the gap between molecular structures and textual descriptions. Despite this intention, LLMs trained on ChEBI-20 have demonstrated limited effectiveness in our TOMG-Bench benchmark. For instance, BioT5-base, which is claimed as the state-of-the-art (SOTA) model for text-based molecule generation on the ChEBI-20 dataset, only achieves a weighted average accuracy of 4.21% on the TOMG-Bench. In the MolEdit and MolOpt tasks, these models are unable to execute correct operations on provided molecules, resulting in disappointing similarity scores. Similarly, in the MolCustom task, which closely mirrors the text-based molecule generation task, the performance remains unsatisfactory, with no model achieving a score above 5% in a single subtask. This performance shortfall highlights a critical limitation of the ChEBI-20 dataset, as it lacks the data quantity and diversity necessary to effectively align molecules with textual descriptions.
In contrast, TOMG-Bench offers a more comprehensive and intricate evaluation framework for text-to-molecule generation. With a larger and more varied set of test examples, TOMG-Bench could robustly assess the capabilities of language models in translating textual descriptions into molecular structures. As such, it represents a significant advancement in the evaluation of text-based molecule generation.
OpenMolIns can enable LLMs to achieve better performance than the most powerful open source general LLMs. We have also developed OpenMolIns, an instruction-tuning dataset, to enhance LLMs’ proficiency in the tasks outlined in the TOMG-Bench. Across five distinct data scales, we observed a pronounced data scaling law: as the size of the corpus increases, the performance of LLMs also improves. In particular, for Galactica-125M, we assessed its capabilities comprehensively on both five distinct data scales. As shown in Figure 3, the outcomes were remarkable: Galactica-125M achieved a weighted average score of 25.73% on OpenMolIns-xlarge, surpassing even the 70B Llama-3-70B-Instruct and GPT-3.5-turbo, with only 125 million parameters. Meanwhile, the results of Galactica-125M show a clear data scaling law, denoting that LLMs are hungry for more molecule corpora to achieve better performance. Notably, OpenMolIns-large has also enabled Llama3.1-8B-Instruct to outperform all the existing open-source general LLMs in TOMG-Bench, showing the effectiveness of the dataset.

### v2

@@ 3.3 Metric Design
Task (Weighted) Success Rate Similarity Novelty Validity
MolEdit
MolCustom
MolOpt
Table 1: All used metrics are listed here. ’✓’ means that a metric is calculated on a task.
3.3.1 MolEdit & MolOpt Evaluation
@@
For the MolEdit and MolOpt tasks, as shown in Table 1, the following metrics are adopted:
Success Rate: We have developed automated testing processes with RDKit, which compare the generated molecule with the original molecule and check if the number of the specified functional group is correct after modification (for MolEdit task) and if the specified chemical property is optimized as required (for MolOpt task). Details of the automated testing process are shown in Appendix E. The success rate reflects the percentage of molecules that have successfully passed the rigorous testing processes. However, as it is impossible to track every change between the generated molecules and the original molecule (See Appendix E for detailed explanation), successfully passing the test processes does not necessarily mean that the generated molecule is modified from the original molecule.
Similarity: In this case, we need to evaluate the similarity between the original molecule and the generated molecule. When removing a specified functional group, the modification should follow the minimal change principle. Therefore, a high similarity indicates that the generated molecule can be modified from the original molecule via as few as possible steps, instead of transforming the molecule into an entirely new one or inadvertently removing additional components. For implementation, we calculate the Tanimoto Similarity between the generated molecule and the original molecule by transforming them into the Morgan Fingerprints (Butina, 1999), which can be described by the following equation:
$\delta(m^{g},m^{o})=\frac{\left|fp_{m^{g}}\cap fp_{m^{o}}\right|}{\left|fp_{m^{g}}\cup fp_{m^{o}}\right|},$ (1)
where $m^{g}$ denotes the generated molecule, $m^{o}$ is the original molecule, and $\delta(m^{g},m^{o})$ is the Tanimoto Similarity between the generated molecule and the original molecule, while $fp_{m^{g}}$ and $fp_{m^{o}}$ represents their corresponding Morgan Fingerprints. $\left|fp_{m^{g}}\cap fp_{m^{o}}\right|$ is the size of the intersection between their Morgan Fingerprints, while $\left|fp_{m^{g}}\cup fp_{m^{o}}\right|$ is the union.
Then, the similarity score for a subtask $t$ in MolEdit and MolOpt, denoted as $\delta_{t}$, is determined by taking the average of the similarity scores from all individual test samples.
Validity: This metric ensures that the generated molecules are chemically valid and follow the rules of molecular syntax. We use RDKit to parse the SMILES strings. Only the molecules that successfully pass the parser are considered valid molecules. A higher validity means that the model is more familiar with the grammar of molecule textual representations.
3.3.2 MolCustom Evaluation
For the MolCustom task, the metrics adopted are also shown in Table 1. As LLMs are required to generate the molecule from scratch, we employ a different testing process for the evaluation of success rate and introduce the novelty score instead of similarity:
Success Rate: For different subtasks in MolCustom, we have designed different testing processes by the chemical toolbox. We calculate the number of the specified atoms, functional groups, or bonds for the three subtasks to check if the generation meets the requirements. Detailed testing process of MolCustom could also be found in Appendix E. A high success rate indicates that the LLM is effectively generating molecules according to the request.
Novelty: However, since there may be numerous valid solutions, fulfilling the basic requirements alone does not guarantee the quality of generation. To enable the exploration of new molecule structures, we employ novelty as an indication of the quality, which compares the similarity between the newly generated molecules and existing ones in the database. The novelty score is particularly important for discovering potential drug or material candidates, where innovation can lead to significant advancements. For implementation, we choose the Zinc250K as the database and calculate the average Tanimoto Similarity between the generated molecules and the existing molecules. The novelty $n$ for the generated molecule $m^{g}$ can be calculated as follow:
$n(m^{g})=1-\frac{\sum_{m^{k}\in Zinc250K}\delta(m^{g},m^{k})}{\left|Zinc250K\right|},$ (2)
Then, the novelty score forthe subtask $t$ in MolCustom, denoted as $n_{t}$, is determined by taking the average of the novelty scores from all individual test samples.
Validity: Similarly, validity is also an important metric for the MolCustom task.
3.3.3 Average Weighted Success Rate
@@
To comprehensively evaluate the average performance of LLMs on TOMG-Bench, we introduce an average weighted success rate to rank the performance of LLMs. Considering that the success rate alone can not ensure the accurate generation for MolEdit and MolOpt tasks and the novelty score for MolCustom is also a critical metric for the quality of molecule generation, we adopt similarity and novelty scores as weights of the success rate:
$W\!S\!R_{t}=\begin{cases}n_{t}\times S\!R_{t},&t\in\{MolCustom\}\\ \delta_{t}\times S\!R_{t},&t\in\{MolEdit,MolOpt\}\end{cases},$ (3)
where $W\!S\!R_{t}$ denotes the weighted success rate for a subtask $t$, while $\delta_{t}$ is the similarity score for the MolEdit and MolOpt tasks, $n_{t}$ represents the novelty score for the MolCustom tasks, and $S\!R_{t}$ is the corresponding success rate of the subtask. Then, the average weighted success rate $\overline{W\!S\!R}$ could evaluate the synthetic performance of LLMs among all the nine subtasks:
$\overline{W\!S\!R}=\frac{1}{9}\sum_{t}W\!S\!R_{t},$ (4)
@@
where the average weighted success rate $\overline{W\!S\!R}$ provides a balanced measure of performance that considers both the success rate and quality of the generation.
@@ 3.4 OpenMolIns: Instruction Tuning Dataset
We collect the instruction tuning dataset by constructing the samples of the nine subtasks in equal amounts. The RDKit toolbox is applied to construct the instruction tuning dataset. Details of the construction process and the pre-defined prompt templates are shown in Appendix D.
For the MolEdit and MolOpt domains, we randomly select a functional group from the original molecule for addition or removal. Then, similarly, we wrap the original molecule and the edited molecule with a random pre-defined prompt to construct the training sample for MolEdit. In the case of MolOpt, we also evaluate the direction of the desired property changes by calculating the properties to determine whether the operation improves or decreases the property value. Then, the original molecule, the optimized molecule, and the optimization direction are altogether wrapped by a pre-defined prompt.
In the MolCustom domain, constructing the instruction tuning samples is rather straightforward. We calculate the molecular statistics and encapsulate them within prompts. For example, for AtomNum, we count all the atoms we would consider in the molecule, including their types and numbers. Then, we wrap these statistics with a random pre-defined prompt to construct the training sample. This method allows the generated molecules to better fit the real distribution of molecular space.
To investigate the impact of data scales on the performance of LLMs, we have established five distinct data levels: light, small, medium, large, and xlarge, as illustrated in Section 3.5. Each level represents a different quantity of data, shown in Table 2, allowing researchers to analyze how the amount of training data influences the model’s ability to generate and edit molecules effectively.
Dataset TOMG-Bench OpenMolIns
item each subtask total light small medium large xlarge
Data Size 5,000 45, 000 4,500 18,000 45,000 90,000 1,200,000
@@ 3.5 Statistics
In this section, we outline the data sizes of TOMG-Bench along with the OpenMolIns dataset, which are shown in Table 2. For TOMG-Bench, we have three main tasks with nine subtasks in total, where each subtask contains 5,000 test samples.
OpenMolIns has five distinct data scales: light, small, medium, large, and xlarge, ranging from 4,500 to 1,200,000 examples, which helps us investigate the data scaling law of applying LLMs to the text-based open molecule generation task. The nine subtasks within the TOMG-Bench are uniformly distributed in OpenMolIns.
@@ 4 Experiments
4.1 Models
@@
Proprietary Models. This category includes LLMs that are only accessible via commercial API services. In this work, we benchmark GPT-4o, GPT-4-turbo, GPT-3.5-turbo (OpenAI, 2023), Claude-3.5 (Anthropic, 2024b), Claude-3 (Anthropic, 2024a), and Gemini-1.5-pro (Deepmind, 2024), which are all the most advanced LLMs with powerful language understanding and generation capabilities. These LLMs are pre-trained on vast pre-training corpora, normally including chemical-related knowledge.
Open-source General LLMs. This group contains open-source LLMs that are tuned with the instruction following capability, which can be used for a wide range of tasks and applications. Specifically, we benchmark Llama-3-70B-Instruct, Llama-3-8B-Instruct, Llama-3.1-8B-Instruct, Llama-3.2-1B-Instruct (Dubey et al., 2024), Mistral-7B-Instruct-v0.2 (Jiang et al., 2023), Qwen2-7B-Instruct (Yang et al., 2024), yi-1.5-9B (Young et al., 2024), and chatglm-9B (GLM et al., 2024). These LLMs are top-tier open source general LLMs within our computation budget.
Open-source ChEBI-20 Fine-tuned LLMs. LLMs fine-tuned on the ChEBI-20 dataset can grasp some extent of text-based molecule generation capability. In this case, our experiments also cover LLMs like MolT5-small, MolT5-base, MolT5-large (Edwards et al., 2022), and BioT5-base (Pei et al., 2023). These models have been the state-of-art models in the molecule-caption translation task.
OpenMolIns Fine-tuned LLMs We further adopt LLMs like Galactica-125M (Taylor et al., 2022), Llama3.2-1B-Instruct, and Llama-3.1-8B-Instruct on OpenMolIns dataset for instruction tuning. We specifically include the experiments on five distinct data sizes of OpenMolIns for Galactica-125M. We select Galatica-125M because it has been pre-trained on scientific corpora and has proved its effectiveness in molecule-related tasks (Liu et al., 2023). Meanwhile, a small size model can also help us study the data scaling law within reasonable budget. Besides, Llama3.2-1B-Instruct and Llama-3.1-8B-Instruct are selected due to their advanced capabilities and similar architectures.
Figure 3: The performance of models benchmarked in TOMG-Bench. In TOMG-Bench, LLMs are divided into 4 categories: Proprietary Models, Open-source General LLMs, Open-source ChEBI-20 Fine-tuned LLMs, and OpenMolIns Fine-tuned LLMs. Models of unknown parameters are denoted as horizontal lines.
@@ 4.2 Findings
Figure 3 presents the performances and model sizes of different models benchmarked on TOMG-Bench, as well as the instruction-tuning performance of Galactica-125M on the five distinct data levels of OpenMolIns. More precise details for each subtask are further illustrated in Appendix F. Based on the above results, we observe the following key findings:
Text-based open molecule generation is both promising and challenging for LLMs. As illustrated in Figure 3, we have calculated the average weighted success rate across all the nine subtasks. Among the LLMs benchmarked, Claude-3.5 stands out as the top performer, achieving an average weighted success rate of 35.92%. Gemini-1.5-pro follows closely with an average weighted success rate of 34.80%. Meanwhile, fine-tuned on OpenMolIns-large, Llama-3.1-8B-Instruct achieves an average weighted success rate of 27.22%, outperforming all the open-source general LLMs.
However, these results still underscore the considerable scope for improvement, even among the most advanced proprietary LLMs. What’s more, despite LLMs like Claude-3.5 and GPT-4o exhibit relatively strong performance in the MolEdit and MolOpt tasks, the more challenging MolCustom task still remains a challenge. In MolCustom tasks, no LLM has managed to achieve a success rate exceeding 25% for a single subtask. Meanwhile, as shown in Figure 5, among the three subtasks of MolCustom, BondNum is the most difficult task, as it requires LLMs to understand the connections or edges in the molecule graph merely from the textual representation. These observations indicate that the generation of molecules from scratch demands a deep understanding of the molecular structures, an area where current models are still striving to make significant strides.
Most powerful open-source general LLMs can already outperform GPT-3.5-turbo. In TOMG-Bench, Llama-3-70B-Instruct has achieved an impressive average weighted success rate of 23.93%, notably outperforming GPT-3.5-turbo, which scored 18.58%. Compared to the proprietary models, open-source general LLMs seem to have been lack of chemical corpora for the pre-training process. However, recent development of open-source general LLMs have rapidly bridged the gap. The evolution of the Llama series, in particular, has been remarkable, finally surpassing GPT-3.5-turbo and showing the increasing diversity of pre-training corpora.
More powerful LLMs usually inherit a better performance in TOMG-Bench. Across all the LLMs we benchmarked, a clear trend emerged: the more powerful the LLM is, the higher performance it can achieve on the TOMG-Bench. For instance, the GPT series has consistently demonstrated improved performance with each new iteration from GPT-3.5-turbo to GPT-4o.
Similarly, within the Llama-3 series, we could also observe that larger and more powerful models tend to achieve superior results on the TOMG-Bench. These findings underscore a strong correlation between an LLM’s capabilities and its performance in text-based open molecule generation.
@@
ChEBI-20 is insufficient for LLMs to master molecular structures and editing operations. The ChEBI-20 dataset is designed to bridge the gap between molecules and textual descriptions. Despite this intention, LLMs trained on ChEBI-20 have demonstrated limited effectiveness in TOMG-Bench. For instance, BioT5-base, which is claimed as the state-of-the-art model for text-based molecule generation on the ChEBI-20 dataset, only achieves an average weighted success rate of 4.21% on the TOMG-Bench. In the MolEdit and MolOpt tasks, although these models obtain a high success rate, the generations are actually not modified from the original molecules, leading to poor similarity scores. This also shows that the average weighted success rate provides a more accurate reflection of the actual performance. Similarly, in the MolCustom task, which closely mirrors the text-based molecule generation task, the performance remains unsatisfactory, with no model achieving a success rate above 5% in a single subtask. This performance shortfall highlights a critical limitation of the ChEBI-20 dataset, as it lacks the data quantity and diversity necessary to effectively align molecules with textual descriptions.
The data scaling law varies among different tasks. Across five distinct data scales of OpenMolIns and the experiments on Galactica-125M, we observe a general data scaling law: as the size of the corpus increases, the performance of LLMs also improves. As shown in Figure 3 and Table 12, the outcomes were remarkable: Galactica-125M achieved a weighted average score of 25.73% on OpenMolIns-xlarge, surpassing even the 70B Llama-3-70B-Instruct and GPT-3.5-turbo, with only 125 million parameters.
However, an interesting observation shows that the data scaling law varies among different tasks. For instance, in MolCustom tasks, we witnessed a substantial average improvement of 286% when we scaling the data size from large to xlarge, which suggests that, for specific tasks, increasing data size has a beneficial effect on model performance.
Conversely, in the MolCustom tasks, the performance gains from data scaling were notably modest. More concerning, in the MolEdit tasks, we even observed a decline in performance with increased data size, which indicates that the model’s capacity is not solely data-driven but is also constrained by the number of parameters. It suggests that for certain subtasks within TOMG-Bench, simply increasing the volume of data may not lead to proportional improvements in performance, and other factors such as model complexity or optimization strategies may need to be considered.